Solve the system of equations. If the system has no solution, say that it is inconsistent.
15x-6y=6
8x+4y=2
X=_____ Y=_____

Answer by stanbon(75887)

Solve the system of equations. If the system has no solution, say that it is inconsistent.
15x-6y=6
8x+4y=2
-----
Simplify::
5x - 2y = 3
4x + 2y = 1
----
9x = 4
x = 4/9
-----
Solve for "y"::
4(4/9) + 2y = 1
2y = 1 - 16/9
2y = -7/9
y = -7/18
--------------
X= 4/9 ; Y= -7/18
--------------
Ans: Consistent
-----------
Cheers,
Stan H.
-----------